# RoboHorizon: An LLM-Assisted Multi-View World Model for Long-Horizon Robotic Manipulation

### v1

@@ 4 Experiments
We design our experiments to explore the following questions: (i) How does RoboHorizon, following the RSPA pipeline, perform compared to SPA-driven model-based reinforcement learning (RL) baselines in short- and long-horizon manipulation tasks? (ii) If SPA-driven baselines also adopt LLM-generated dense reward structures, does RoboHorizon still maintain its advantage? (iii) To what extent do RoboHorizon’s key design components impact its overall performance?
@@ 4.1 Performance Comparison
SPA-driven baselines with LLM-generated dense rewards vs. RoboHorizon
To compare the performance of SPA-driven baselines with LLM-generated dense rewards and RoboHorizon, we replace the manually defined reward system in SPA-driven baselines with the dense reward mechanism generated during RoboHorizon’s Recognize phase using LLM assistance. Fig. 5 shows the success rate comparison between RoboHorizon and five SPA-driven baselines enhanced with LLM-generated dense rewards (as indicated by “LLM” in the legend) across 4 short-horizon tasks and 9 long-horizon tasks. It also illustrates the performance improvement of SPA-driven baselines with LLM-generated dense rewards compared to manually defined rewards. The results show that even with the introduction of LLM-generated dense reward structures, RoboHorizon consistently demonstrates superior performance over all baselines. However, these baselines achieve significant performance improvements when using LLM-generated dense rewards, particularly in long-horizon tasks, which highlights the effectiveness of LLM-generated reward mechanisms in helping robots comprehend long-horizon tasks. Nevertheless, due to the limitations of SPA-driven baselines in handling complex long-term task representations, even with LLM-generated rewards, their performance in long-horizon tasks remains inferior to RoboHorizon.
This result clearly answers our second question: with LLM-generated rewards, SPA-driven baselines improve their performance in both short-horizon and long-horizon tasks, especially in long-horizon tasks. However, due to their limited ability to learn complex long-horizon representations, RoboHorizon still holds a significant performance advantage. This finding further underscores the importance of the LLM-assisted reward generation mechanism and the key-horizon multi-view representation learning modules. These modules not only enhance the robot’s ability to interpret task instructions but also improve its perception of tasks and target objects in complex long-horizon scenarios.

### v2

@@ 4 Experiments
We design our experiments to explore the following questions: (i) How does RoboHorizon, following the RSPA pipeline, perform compared to SPA-driven model-based reinforcement learning (RL) baselines in short- and long-horizon manipulation tasks? (ii) If SPA-based baseline methods also adopt stepwise rewards generated by LLMs but do not aware the staged reward for each sub-task stage, does RoboHorizon still maintain its advantage? (iii) To what extent do RoboHorizon’s key design components impact its overall performance?
@@ 4.1 Performance Comparison
SPA-driven baselines with LLM-generated stepwise rewards vs. RoboHorizon
To compare the performance of SPA-driven baselines with LLM-generated stepwise rewards but without staged rewards for each sub-task phase against RoboHorizon, we replace the manually defined reward system in SPA-driven baselines with the stepwise rewards generated during RoboHorizon’s Recognize phase using LLM assistance. Fig. 5 shows the success rate comparison between RoboHorizon and five SPA-driven baselines enhanced with LLM-generated stepwise rewards (as indicated by “LLM” in the legend) across 4 short-horizon tasks and 9 long-horizon tasks. It also illustrates the performance improvement of SPA-driven baselines with LLM-generated stepwise rewards compared to manually defined rewards. The results show that while these baseline methods achieve some improvement in task success rates when using LLM-generated stepwise rewards, demonstrating the effectiveness of well-designed stepwise rewards aligned with motion planning in helping robots understand operational tasks, their performance gains in long-horizon tasks remain very limited without considering staged rewards for each sub-task phase. Furthermore, regardless of the task, they fail to surpass the full RoboHorizon framework driven by our proposed RSPA pipeline.
This result clearly answers our second question: By using LLM-generated stepwise rewards aligned with motion planning, SPA-driven baseline methods achieve certain performance improvements in both short-horizon and long-horizon tasks. However, due to their lack of consideration for staged rewards in each sub-task phase and their limited ability to learn complex long-horizon task representations, RoboHorizon still maintains a significant performance advantage. This finding further highlights the importance of the LLM-assisted reward generation mechanism and key-horizon multi-view representation learning modules in the RSPA-driven RoboHorizon framework. These modules not only enhance the robot’s ability to interpret task instructions but also improve its perception of tasks and target objects in complex long-horizon scenarios.